NEED ANSWER ASAP / ANSWER NEVER USED BEFORE

a.)

Forecasted Statements and Ratios

Upton Computers makes bulk purchases of small computers, stocks them in conveniently located warehouses, ships them to its chain of retail stores, and has a staff to advise customers and help them set up their new computers. Upton's balance sheet as of December 31, 2019, is shown here (millions of dollars):

Cash $   3.5 Accounts payable $   9.0
Receivables 26.0 Notes payable 18.0
Inventories 58.0 Line of credit 0
Total current assets $ 87.5 Accruals 8.5
Net fixed assets 35.0 Total current liabilities $ 35.5
Mortgage loan 6.0
Common stock 15.0
Retained earnings 66.0
Total assets $122.5 Total liabilities and equity $122.5

Sales for 2019 were $500 million and net income for the year was $15 million, so the firm's profit margin was 3.0%. Upton paid dividends of $6 million to common stockholders, so its payout ratio was 40%. Its tax rate was 25%, and it operated at full capacity. Assume that all assets/sales ratios, (spontaneous liabilities)/sales ratios, the profit margin, and the payout ratio remain constant in 2020. Do not round intermediate calculations.

  1. If sales are projected to increase by $110 million, or 22%, during 2020, use the AFN equation to determine Upton's projected external capital requirements. Enter your answers in millions. For example, an answer of $10,550,000 should be entered as 10.55. Round your answer to two decimal places.

    $    million

  2. Using the AFN equation, determine Upton's self-supporting growth rate. That is, what is the maximum growth rate the firm can achieve without having to employ nonspontaneous external funds? Round your answer to two decimal places.

      %

  3. Use the forecasted financial statement method to forecast Upton's balance sheet for December 31, 2020. Assume that all additional external capital is raised as a line of credit at the end of the year and is reflected (because the debt is added at the end of the year, there will be no additional interest expense due to the new debt). Assume Upton's profit margin and dividend payout ratio will be the same in 2020 as they were in 2019. What is the amount of the line of credit reported on the 2020 forecasted balance sheets? (Hint: You don't need to forecast the income statements because the line of credit is taken out on the last day of the year and you are given the projected sales, profit margin, and dividend payout ratio; these figures allow you to calculate the 2020 addition to retained earnings for the balance sheet without actually constructing a full income statement.) Enter your answers in millions. For example, an answer of $10,550,000 should be entered as 10.55. Round your answers to two decimal places.

    Upton Computers
    Pro Forma Balance Sheet
    December 31, 2020
    (Millions of Dollars)
    Cash $   
    Receivables $   
    Inventories $   
    Total current assets $   
    Net fixed assets $   
    Total assets $   
    Accounts payable $   
    Notes payable $   
    Line of credit $   
    Accruals $   
      Total current liabilities $   
    Mortgage loan $   
    Common stock $   
    Retained earnings $   
      Total liabilities and equity $   

b.)

Projected Operating Assets

Berman & Jaccor Corporation's current sales and partial balance sheet are shown below.

This year
Sales $ 1,000
Balance Sheet: Assets
Cash $ 150
Short-term investments $ 120
Accounts receivable $ 150
Inventories $ 200
    Total current assets $ 620
Net fixed assets $ 450
    Total assets $ 1,070

Sales are expected to grow by 8% next year. Assuming no change in operations from this year to next year, what are the projected total operating assets? Do not round intermediate calculations. Round your answer to the nearest dollar.

$  
Answer #1
a] UPTON:
1] AFN is given by the equation:
AFN = (A/S) x (Δ Sales) - (L/S) x (Δ Sales) - (PM x FS x (1-d))
where,
A / S = Assets that change givendirectly with sales.
Δ sales = Change in sales
L / S = Liabilities that change directly with sales
PM = Profit Margin on Sales = net income / sales.
FS = Forecasted Sales
d = dividend payout ratio
(1 - d) = retention ratio
Substituting required values in the above equation we have,
AFN = (122.5/500)*110-((9+8.5)/500))*110-610*3%*(1-40%) = $             12.12
2] Setting AFN = 0
0 = (122.5/500)*x-(17.5/500)*x-(500+x)*0.018, where x is the
dollar increase in sales.
Solving for x
0 = 0.245*x-0.035*x-9-0.018*x
9 = 0.192*x
x = 9/0.192 = $          46.875
Self supporting growth rate = 46.875/500 = 9.38%
3] Pro Forma Balance Sheet
December 31, 2020
(Millions of Dollars)
Cash = 3.5*122% = $               4.27
Receivables = 26*122%= $             31.72
Inventories = 58*122% = $             70.76
Total current assets $          106.75
Net fixed assets = 35*122% = $             42.70
Total assets $          149.45
Accounts payable = 9*122% = $             10.98
Notes payable $             18.00
Line of credit [51.47-10.98-18-10.37] $             12.12
Accruals = 8.5*122% = $             10.37
  Total current liabilities [149.45-76.98-15-6] $             51.47
Mortgage loan $               6.00
Common stock $             15.00
Retained earnings [66+610*1.8%] $             76.98
  Total liabilities and equity $          149.45
b] BERMAN & JACCOR CORPORATION:
Projected total operating assets = Current total operating assets*(1+growth rate) = (1070-120)*(1+0.08) = $       1,026.00
Note: Short term investments do not form part of operating
assets.